Consider an elementary aqueous reversible reaction system, A↔︎B. The equilibrium constant at 62°C is 11.5. What is the equilibrium conversions at temperature 62°C ?______________

1
Expert's answer
2022-02-15T03:37:02-0500

KT = Kc(RT)

Kc = KT/(RT) = 11.5 / (8.31 x (62+273)) = 0.004